# Multi-Agent Reinforcement Learning with Control-Theoretic Safety Guarantees for Dynamic Network Bridging

### v1

@@ Abstract
Addressing complex cooperative tasks in safety-critical environments poses significant challenges for Multi-Agent Systems, especially under conditions of partial observability. This work introduces a hybrid approach that integrates Multi-Agent Reinforcement Learning with control-theoretic methods to ensure safe and efficient distributed strategies. Our contributions include a novel setpoint update algorithm that dynamically adjusts agents’ positions to preserve safety conditions without compromising the mission’s objectives. Through experimental validation, we demonstrate significant advantages over conventional MARL strategies, achieving comparable task performance with zero safety violations. Our findings indicate that integrating safe control with learning approaches not only enhances safety compliance but also achieves good performance in mission objectives.
@@ I Introduction
This challenge highlights a fundamental limitation of all Reinfocement Learning (RL)-based approaches: while agents learn to optimize predefined reward functions, their behavior lacks guarantees and remains inherently unpredictable. The sole reliance on MARL’s reward functions to ensure safety is insufficient [1]. Even with extensive training and monitoring of an agent’s performance based on expected rewards, or other performance measures, it is impossible to exhaustively examine all potential scenarios in which the agent might fail to act safely or predictably.
In ensuring safety among agents, various control-theoretic approaches and hybrid learning approaches have been developed, ranging from Model Predictive Control (MPC) [2] to Reference Governor (RG) techniques [3], and from Control Barrier Functions (CBFs) [4] to Reachability Analysis (RA) methods [5]. In this work, we focus on developing a hybrid learning approach built upon the method used in [6], that can learn to achieve complex objectives while ensuring safety guarantees in distributed MAS.
To demonstrate and evaluate the effectiveness of our hybrid approach we consider the cooperative task of Dynamic Network Bridging [7] (Fig. 1). In this task the goal is to establish and maintain a connection between two moving targets A and B in a 2D plane, relying on a swarm of $N$ agents with limited observation of the environment. The agents must form an ad-hoc mobile network that establishes a communication path between the two targets as they move, dynamically adjusting their positions to ensure an uninterrupted multi-hop communication link.
@@
In this paper we introduce a hybrid approach combining MARL and control-theoretic methods, accomplishing task objectives while providing safety guarantees. Our key contributions are: 1) A decentralized control framework that is MARL compatible and restricts the effect of each agent’s movement updates to only its one-hop neighbors, enabling efficient local coordination; 2) An algorithm that updates setpoints while preserving safety conditions through communication with affected neighbors; 3) An analytical computationally tractable condition verifying potential safety violations during updates. Overall our approach provides effective coordination among agents while ensuring safety constraints.
@@ II Preliminaries
In this section, we present the foundational theoretical elements, drawing from both control theory and MARL, that form the basis of our proposed hybrid approach.
Consider a linear time invariant (LTI) system expressed in the state space form
$\dot{x}=Ax+Bu$ (1)
where $x\in\mathbb{R}^{n}$ is the state vector with $n$ components, and $u\in\mathbb{R}^{p}$ is the input vector with $p$ components. Thus, the dynamical matrices are $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times p}$. If the pair $(A,B)$ is controllable, there exists a state feedback controller:
$u=-Kx$ (2)
such that, the origin ($x=0$) of the resulting closed-loop (C-L) system:
$\dot{x}=(A-BK)x=A_{f}x$ (3)
is an asymptotically stable equilibrium point (EP): $x(t)\rightarrow 0\quad\text{as}\quad t\rightarrow+\infty$. Hence, the matrix $A_{f}$ is Hurwitz, i.e. all the eigenvalues have a negative real part [8]. The problem to design (2) that makes (1) asymptotically stable is called stabilization problem.
II-A Tracking Control Problem
In the case the problem requires to move the system (1) following a specific trajectory, the problem is called tracking control. This problem can be solved by generating a sequence of EPs [9] that we call setpoints, denoted by $x_{sp}$. The C-L system (3) can be represented as:
$\dot{x}=A_{f}(x-x_{sp})$ (4)
where the control input (2) becomes $u=-K(x-x_{sp}).$
II-B Lyapunov’s Theory
Since $A_{f}$ is Hurwitz, there exists a symmetric positive definite matrix $P>0$ such that:
$A_{f}^{T}P+PA_{f}=-Q$ (5)
for a given symmetric positive definite matrix $Q>0$. Equation (5) is called Lyapunov’s equation. Defining the $P$-norm as:
$\lVert x-x_{sp}\rVert_{P}\triangleq\sqrt{(x-x_{sp})^{T}P(x-x_{sp})}$ (6)
the $P$-norm squared is a Lyapunov function $V(x)=\lVert x-x_{sp}\rVert_{P}^{2}$, which is always positive except at $x_{sp}$ where it is zero. For all trajectories $x(t)$ starting within the ellipsoid:
$\mathcal{E}_{c}(x_{sp})=\{x\in\mathbb{R}^{n}\,|\,\lVert x-x_{sp}\rVert_{P}^{2}\leq c\},$ (7)
where $c>0$, $V(x(t))$ is monotonically decreasing. Therefore, the trajectory $x(t)$ never leaves $\mathcal{E}_{c}$, which is called a positively invariant set [8].
Remark 1
In the case of a nonlinear dynamical system described by $\dot{x}=f(x,u)$ where $u=0$, $x=0$ is an equilibrium point for the system $f(0,0)=0$. The control input (2) can be designed by considering the linearized system (1) around the equilibrium point. Therefore, the closed-loop (CL) system (3) can be used to solve Lyapunov’s equation (5), which is used to compute the positively invariant set $\mathcal{E}_{c}(x_{sp})$ (7). While for LTI systems, $\mathcal{E}_{c}(x_{sp})$ is positively invariant for any positive $c\in\mathbb{R}$, for nonlinear systems, we need to verify that condition in a neighborhood of the equilibrium point $x_{sp}$ [8]. Furthermore, for nonlinear systems, $x_{sp}$ has to satisfy $f(x_{sp},0)=0$.
II-C 6-DOF UAV
In this work, we consider a 6-DOF UAV as the agent, described by a state vector of 12 components:
$x=\begin{bmatrix}\dot{p}_{x},&p_{x},&\dot{p}_{y},&p_{y},&\dot{p}_{z},&p_{z},&\dot{\phi},&\phi,&\dot{\theta},\\ \theta,&\dot{\psi},&\psi\end{bmatrix}^{T}$ (8)
where $p_{x}$, $p_{y}$, $p_{z}$, and their derivatives represent linear position and velocity, respectively. Similarly, $\phi$, $\theta$, $\psi$, and their derivatives represent the Euler angles (roll, pitch, and yaw) and their angular velocities, respectively. This model is consistent with the one used in [9] and references therein. The controller used is given by (2), obtained by linearizing the quadrotor model around an equilibrium point. For this case, the equilibrium point is indicated with $x_{sp}$ which has the same structure of (8) with all components equal to zero except for $p_{x}$, $p_{y}$, $p_{z}$.
II-D Multi-Agent Reinforcement Learning
@@
A Dec-POMDP is defined by a tuple $(S,\mathbf{A},\mathcal{O},T,R,\Omega,\gamma,N)$ where $S$ is a finite set of states of the environment, $\mathbf{A}$ is a set of joint action spaces, and $N$ is the number of agents, $T$ is the state transition probability function which gives the probability of transitioning from state $s$ to state $s^{\prime}$ when joint action $\mathbf{a}=(a_{1},\cdots,a_{N})$ is taken; $R$ is the joint reward function that maps states and joint actions to real numbers and is used to specify the goal of the agents; $\Omega$ is a set of joint observations $O_{i}$, where $O_{i}$ is the observation set for agent $i$; $\mathcal{O}$ is the joint observation probability function, which gives the probability of receiving joint observation $\mathbf{o}=(o_{1},\cdots,o_{N})$ after taking joint action $\mathbf{a}$ and ending up in state $s^{\prime}$; and $\gamma$ is the reward discount factor, $0\leq\gamma\leq 1$.
In a Dec-POMDP, agents need to collaborate and coordinate their actions based on their individual observations to maximize the joint reward function. Solving a Dec-POMDP optimally is computationally intractable, as the policy space grows exponentially with the number of agents and the planning horizon. Therefore, various approximate solution methods [12, 13] have been used, such as heuristic search, dynamic programming, value or policy-based approximators with neural networks etc. The problem facing the team is to find the optimal joint policy, i.e. a combination of individual agent policies that produces behavior that maximizes the team’s expected reward.
III Problem Formulation
We address the challenge of dynamically forming and sustaining a connection between two moving targets, A and B, in a 2D space using a swarm of $N$ agents [7]. These agents must collaboratively adjust their positions to maintain a continuous multi-hop link between the targets, operating under decentralized control and limited local perception, while ensuring they avoid collisions and respect physical boundaries. We consider all the agents as nonlinear systems that can be reduced into the form (1) controlled by (2) around the equilibrium point $x_{sp,i}$ [9]:
$\dot{x}_{i}=A_{f}(x_{i}-x_{sp,i})=(A-BK)(x_{i}-x_{sp,i})$ (9)
for $i=1,\ldots,N$. In this case, we consider $N$ agents described with the same model, but in general, our approach can be generalized to the case of different dynamics and multidimensional scenarios.
III-A One-hop Communication
We assume that for each agent, there is a ball of radius $r$ in the 2- or 3-dimensional Euclidean space centered at the agent’s current linear position $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t),p_{z,i}(t)]^{T}$:
Given two agents $i$ and $j$ with $i\neq j$, they can communicate if and only if
$\mathcal{B}_{r}(p_{i}(t))\cap\mathcal{B}_{r}(p_{j}(t))\neq\emptyset$ (10)
In our setup, since $p_{z,i}$ is fixed, the one-hop communication region is described by a 2-D ball. Note that, we use a different notation when we refer to the 2-D communication space to avoid any confusion between the quadrotor $x$-axis and the state space $x$.
III-B Safety
@@
Given two agents $i$ and $j$, where $x_{sp,i}\neq x_{sp,j}$, $x_{i}(t_{0})\in\mathcal{E}_{c}(x_{sp,i})$, and $x_{j}(t_{0})\in\mathcal{E}_{c}(x_{sp,j})$. Agents $i$ and $j$ are safe with respect to each other if
@@
The above definition is derived by the positively invariant property of (7), since the initial conditions of each agent are inside their respective ellipsoids, then $x_{i}(t)\in\mathcal{E}_{c}(x_{sp,i})$ and $x_{j}(t)\in\mathcal{E}_{c}(x_{sp,j})$ for $t\rightarrow+\infty$. This means that the trajectory of each agent cannot intersect generating the risk of possible collisions.
In order to guarantee safety in our setup, we need the following assumption:
@@
The projection of $\mathcal{E}_{c}(x_{sp,i})$ into the 2-D space formed by the components $p_{x}$ and $p_{y}$ must be contained within the 2-dimensional ball $\mathcal{B}_{r}(p_{i}(t))$, where $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t)]$ represents the actual position coordinates ($p_{x,i}(t),p_{y,i}(t)$) of agent $i$ for any state $x_{i}\in\mathcal{E}_{c}(x_{sp,i})$.
III-C Cooperative MARL for Dynamic Network Bridging
Each agent $i$ has a local observation $o_{i}$ comprising the structure of its neighborhood and features representing its neighbors at certain time-step $t$. Such features include the node ID, current coordinates, action taken, and coordinates of the targets A and B to be connected. At the beginning of every time-step, agents make decentralized decisions, generating target points $w$, utilizing their local observation to choose the direction of their next movement. This is encoded in a two-dimensional action space, with each dimension having three options: move forward, move backward, or hold along the corresponding $p_{x}$ or $p_{y}$ axis. Given the direction chosen by the agents, a new target point is calculated for each agent using a fixed offset equal for both the $p_{x}$ and $p_{y}$ axes. Finally, agents are rewarded using a reward function that motivates agents to form a connected network bridging the targets. Such reward signal combines three components:
$R_{\text{base}}(s)=\frac{|C_{\text{max}}(s)|}{|\mathcal{V}|},$ (12)
The overall reward combines these three elements:
We also include a physics simulator, modeling the continuous non-linear dynamics of the agents (2), and a safe tracking control system proposed in the next section (Algorithm 1).
IV Safe Tracking Control
The safety tracking control algorithm that we propose here consists of two steps, the first one is the setpoint update of the agent $i$ that guarantees that its current position belongs to $\mathcal{E}_{c}$ of the new setpoint. This condition is fundamental for the second step where the safety condition (11) is checked for all agents within the one-hop communication ball.
@@ IV Safe Tracking Control
Assuming that the agent $i$ is moving from the target point $w_{i}$ to the target point $w_{i+1}$, generated by the MARL, the idea is to generate a sequence of setpoints on the line that connects the two target points. Once the state of agent $i$, $x_{i}$ reached the setpoint $x_{sp,i}$, a new setpoint $x^{\prime}_{sp,i}$ is generated following the rule
@@
Fig. 2 illustrates the process of updating the setpoint, presented in 2-D space for clearer visualization, though it actually corresponds to a state space with n=12 components, as applicable to quadrotors.
Figure 2: Setpoint update for agent $i$ in the state space. Note that the target points $w_{i}$ and $w_{i+1}$ have the same structure of the setpoints $x_{sp,i}$.
@@
From Assumption 1, the ball $\mathcal{B}_{r}(p_{i}(t))$ encompasses the projection onto the 2-D space of $\mathcal{E}_{c}(x{sp,i})$ for any current state $x_{i}(t)\in\mathcal{E}_{c}(x_{sp,i})$. Utilizing (14) to adjust the setpoint ensures that $x_{i}(t)$ remains within $\mathcal{E}_{c}(x^{\prime}_{sp,i})$, signifying that its 2-D space projection still falls within $\mathcal{B}_{r}(p_{i}(t))$. Thus, should the updated setpoint $x^{\prime}_{sp,i}$ contravene (11), such a violation would only occur concerning agents situated within the one-hop communication range. ∎
Proposition 1 highlights a critical aspect of setpoint updates: setpoint updates only impact safety for agents within one-hop range. This ensures agent $i$ can communicate with affected neighbors before executing updates. This allows coordinating to preemptively address any potential safety concerns from the new setpoint through direct communication with impacted agents.
@@
safety_violation $\leftarrow$ 0;
@@
                   safety_violation $\leftarrow 1$;
@@
       if safety_violation ==absent$=$ = 00 then
@@
Considering the multi-agent system described by (9) and Assumptions 1, 2 and 3, then applying Algorithm 1 to each agent guarantees that the safety condition (11) is satisfied for all agents.
@@
We now need to verify when the safety condition (11) is violated. Our objective is to demonstrate that if (11) is violated, then there exists a point on the line connecting $x_{sp,i}$ and $x_{sp,j}$, inside the overlap region of two ellipsoids $\mathcal{E}_{c}(x_{sp,i})$ and $\mathcal{E}_{c}(x_{sp,j})$ (Fig. 3). This overlap region, denoted by $\mathcal{E}_{c}(x{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})$, satisfies the following condition:
@@
Proposition 3 shows that if there is an ellipsoid violation, then the point $m_{\lambda}$ belongs to the intersection Fig. 3:
@@
Figure 3: Explaination of Proposition 3: $m_{\lambda}$ is on the segment that connects the two setpoints.
@@
Although Algorithm 1 ensures safety according to Proposition 2, it does not guarantee target achievement. Agents within one-hop communication range may become immobilized at a certain position because any attempt to reach the target points could potentially violate the safety condition (11). We intentionally refrain from employing any enforcer to resolve this deadlock scenario, as our aim is to assess whether our MARL framework can autonomously generate target points that circumvent such situations.

### v2

# Safety-Aware Multi-Agent Learning for Dynamic Network Bridging

@@ Abstract
Addressing complex cooperative tasks in safety-critical environments poses significant challenges for multi-agent systems, especially under conditions of partial observability. We focus on a dynamic network bridging task, where agents must learn to maintain a communication path between two moving targets. To ensure safety during training and deployment, we integrate a control-theoretic safety filter that enforces collision avoidance through local setpoint updates. We develop and evaluate multi-agent reinforcement learning safety-informed message passing, showing that encoding safety filter activations as edge-level features improves coordination. The results suggest that local safety enforcement and decentralized learning can be effectively combined in distributed multi-agent tasks.
@@ I Introduction
This challenge highlights a fundamental limitation of all reinforcement learning (RL)-based approaches: while agents learn to optimize predefined reward functions, their behavior lacks guarantees and remains inherently unpredictable. The sole reliance on MARL’s reward functions to ensure safety is insufficient [1]. Even with extensive training and monitoring of an agent’s performance based on expected rewards, or other performance measures, it is impossible to exhaustively examine all potential scenarios in which the agent might fail to act safely or predictably.
In ensuring safety among agents, various control-theoretic approaches and hybrid learning approaches have been developed, ranging from Model Predictive Control (MPC) [2] to Reference Governor (RG) techniques [3], and from Control Barrier Functions (CBFs) [4] to Reachability Analysis (RA) methods [5]. In robotics applications, those methods are known as safety filters [6], and they are deployed to ensure safety regardless of the algorithm used for the agent to accomplish a task. While safety filters can effectively prevent unsafe behavior, they are often conservative by design. This conservatism may lead to conflicts with task objectives, especially in MAS where coordination is essential.
In this work, we develop a training framework for MARL that incorporates safety constraints during policy learning. We adapt an optimization-free safety tracking control algorithm developed in [7], which is based on ellipsoidal positively invariant sets. We show how this safety filter can be adapted to the context of MAS and how safety is guaranteed. We integrate this type of filter in the agents’ training process, allowing them to learn decentralized policies and avoid situations where the filter compromises task objectives.
@@
We demonstrate our approach in the dynamic network bridging task [8] (Fig. 1) where a number of agents establish and maintain a communication path between two moving targets. Our key contributions are as follows: 1) A novel decentralized control framework that integrates a safety filter based on invariant sets into the MARL agents’ training process, and restricts the effect of each agent’s movement updates to only its one-hop neighbors, enabling efficient local coordination; 2) We augment the learning process with safety-informed message passing, where agents observe safety filter activations as edge-level features during training, leading to improved coordination 3) An algorithm that updates setpoints while preserving safety conditions through communication with affected neighbors; 3) An analytical computationally tractable condition verifying potential safety violations during updates. Overall our approach provides effective coordination among agents while ensuring safety constraints.
@@ II Preliminaries
II-A Task description
Dynamic network bridging is a multi-agent task focused on addressing the problem of dynamically establishing and maintaining a communication link between two mobile, independent nodes (targets) using a swarm of autonomous agents taking decentralized decisions [8]. Agents and targets are moving in a 2D grid space and have a limited communication range, within which they exchange information. This region can be defined as a 2D ball of radius $r$, centered at its position $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t),p_{z,i}(t)]^{T}$. Based on this definition, two agents $i$ and $j$ can communicate if and only if their regions intersect:
$\mathcal{B}_{r}(p_{i}(t))\cap\mathcal{B}_{r}(p_{j}(t))\neq\emptyset.$ (1)
where
In our setup, we consider $p_{z,i}$ fixed, therefore the one-hop communication region can be described by a 2D ball. Formally, the system is represented as a dynamic graph $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t))$ with nodes $\mathcal{V}$ corresponding to agents and targets. An edge $(u,v)$ exists between nodes if their Euclidean distance at time-step $t$ is within a predefined communication range.
The task reflects scenarios where communication infrastructure is absent or unreliable, such as in disaster zones or remote environments. Agents must position themselves to form an ad-hoc, decentralized communication network that connects the two mobile targets. Each agent operates based on local observations and limited-range communication, with no access to global state information.
II-B Multi-agent reinforcement learning
@@
A Dec-POMDP is defined by a tuple $(S,\mathbf{A},\mathcal{O},T,R,\Omega,\gamma,N)$ where $S$ is a finite set of states of the environment, $\mathbf{A}$ is a set of joint action spaces, and $N$ is the number of agents, $T$ is the state transition probability function which gives the probability of transitioning from state $s$ to state $s^{\prime}$ when joint action $\mathbf{a}=(a_{1},\cdots,a_{N})$ is taken; $R$ is the joint reward function that maps states and joint actions to real numbers and is used to specify the goal of the agents; $\Omega$ is a set of joint observations $O_{i}$, where $O_{i}$ is the observation set for agent $i$; $\mathcal{O}$ is the joint observation probability function, which gives the probability of receiving joint observation $\mathbf{o}=(o_{1},\cdots,o_{N})$ after taking joint action $\mathbf{a}$ and ending up in state $s^{\prime}$; and $\gamma$ is the reward discount factor, $0\leq\gamma\leq 1$. The goal of the agents is to maximize a joint reward function. Solving a Dec-POMDP optimally is computationally intractable, therefore, various approximate solution methods [11, 12] have been used, such as heuristic search, dynamic programming, value or policy-based approximators with neural networks etc.
II-C Distributed autonomous swarm formation for dynamic network bridging
We formulate our problem as a Dec-POMDP, where each agent has restricted sensing capabilities given by a specific communication range, leading to incomplete information about the global state of the system outside such area. However, whenever two entities (agents or targets) are within range, a communication channel becomes available for them to exchange information. This channel enables the possibility for agents to know each other features and execute distributed strategies requiring message exchange to coordinate. At the beginning of every time-step, agents make decentralized decisions, generating target points $w$, utilizing their local observation to choose the direction of their next movement. To solve the Dec-POMDP that corresponds to our task we used a graph neural network (GNN) approximator that encodes the graph structure of the problem.
II-D Agent Dynamics
We consider a linear time-invariant (LTI) of the form
$\dot{x}=A(x-x_{sp})$ (2)
where $x\in\mathbb{R}^{n}$ is the state of the system, and the origin is translated to the setpoint $x_{sp}\in\mathbb{R}^{n}$. The matrix $A\in\mathbb{R}^{n\times n}$ is Hurwitz (i.e., $\mathrm{Re}(\lambda_{i})<0$), where $\lambda_{i}\in\mathbb{C}$ is the $i$-th eigenvalue of $A$. The matrix $A$ can be seen as the closed-loop system matrix obtained by applying a state-feedback controller (i.e., $u=-K(x-x_{sp})$), where $u$ is the control input.
For the asymptotically stable LTI system (2), there exists a quadratic Lyapunov function of the form
$V(x)=(x-x_{sp})^{T}P(x-x_{sp})$ (3)
where $P$ is a symmetric positive definite matrix that satisfies the Lyapunov equation
$A^{T}P+PA=-Q$ (4)
for a given symmetric positive definite matrix $Q>0$.
Defining the $P$-norm as
$\lVert x-x_{sp}\rVert_{P}\triangleq\sqrt{(x-x_{sp})^{T}P(x-x_{sp})},$ (5)
for a fixed value $c>0$, the Lyapunov function (3) defines an ellipsoid, which can be written in terms of (5) as follows:
$\mathcal{E}_{c}(x_{sp})=\{x\in\mathbb{R}^{n}\mid\lVert x-x_{sp}\rVert_{P}^{2}\leq c\}.$ (6)
From (4), $V(x(t))$ is a monotonically decreasing function; therefore, for any $x(t)$ starting within the ellipsoid (6), the trajectory never leaves it. Hence, $\mathcal{E}_{c}$ is called a positively invariant set [13]. For the rest of the paper, we will refer to the state of the agent $i$ as $x_{i}$ and for the related setpoint as $x_{sp,i}$. In this context, the tracking algorithm updates the setpoint of the agent to reach a desired position. This formulation can be extended to more general nonlinear systems by using linearization [13], or Lyapunov-like methods such as quadratic boundedness [14], as used in the context of UAVs in [7], where ellipsoidal positively invariant sets are provided. We use this last case in our experiments.
II-E Safety
@@
Given two agents $i$ and $j$, with distinct equilibrium points $x_{sp,i}\neq x_{sp,j}$ and initial states $x_{i}(t_{0})\in\mathcal{E}_{c}(x_{sp,i})$ and $x_{j}(t_{0})\in\mathcal{E}_{c}(x_{sp,j})$, respectively, agents $i$ and $j$ are deemed safe relative to each other if:
@@
This definition leverages the invariant properties of (6), ensuring agents’ trajectories remain confined within their respective safety ellipsoids, thereby preventing collisions.
To operationalize this safety criterion within the two-dimensional communication space, we impose the following:
@@
The projection of $\mathcal{E}_{c}(x_{sp,i})$ onto the 2-D plane defined by the coordinates $(p_{x},p_{y})$ must always lie within the 2-D communication ball $\mathcal{B}_{r}(p_{i}(t))$, where $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t)]$ represents the actual position of agent $i$.
III Safety Filter for Tracking Control
The proposed algorithm consists of two steps: first, computing a feasible safe step (Figure 2(a)); and second, checking whether the new position would interfere with another agent (Figure 2(b)).
@@ III Safety Filter for Tracking Control
Assuming that the agent $i$ is moving from the target point $w_{i}$ to the target point $w_{i+1}$, generated by its decentralized policy, the idea is to generate a sequence of setpoints on the line that connects the two target points. Once the state of agent $i$, $x_{i}$ reached the setpoint $x_{sp,i}$, a new setpoint $x^{\prime}_{sp,i}$ is generated following the rule
@@
Fig. 2(a) illustrates the process of updating the setpoint, presented in 2-D space for clearer visualization, though it actually corresponds to a state space with n=12 components, as applicable to quadrotors.
(a) Setpoint update for agent $i$ in the state space.
(b) $m_{\lambda}$ is on the segment that connects the two setpoints.
Figure 2: Safe tracking control and safety verification.
@@
From Assumption 1, the ball $\mathcal{B}_{r}(p_{i}(t))$ encompasses the projection onto the 2-D space of $\mathcal{E}_{c}(x_{sp,i})$ for any current state $x_{i}(t)\in\mathcal{E}_{c}(x_{sp,i})$. Utilizing (8) to adjust the setpoint ensures that $x_{i}(t)$ remains within $\mathcal{E}_{c}(x^{\prime}_{sp,i})$, signifying that its 2-D space projection still falls within $\mathcal{B}_{r}(p_{i}(t))$. Thus, should the updated setpoint $x^{\prime}_{sp,i}$ contravene (7), such a violation would only occur concerning agents situated within the one-hop communication range. ∎
Proposition 1 highlights a critical aspect of setpoint updates: setpoint updates only impact safety for agents within communication range. This ensures agent $i$ can exchange messages with affected neighbors before executing updates. This allows coordinating to preemptively address any potential safety concerns from the new setpoint through direct communication with impacted agents.
@@
potential_safety_violation $\leftarrow$ 0;
@@
                   potential_safety_violation $\leftarrow 1$;
@@
       if potential_safety_violation ==absent$=$ = 00 then
@@
Considering the MAS described where the agents are LTI systems (2) and Assumptions 1, 2 and 3, then applying Algorithm 1 to each agent guarantees that the safety condition (7) is satisfied for all agents.
@@
In the second step, we need to verify whether the safety condition (7) is going to be violated. Our objective is to demonstrate that if (7) is violated, then there exists a point on the line connecting $x_{sp,i}$ and $x_{sp,j}$, inside the overlap region of two ellipsoids $\mathcal{E}_{c}(x_{sp,i})$ and $\mathcal{E}_{c}(x_{sp,j})$ (Fig. 2(b)). This overlap region, denoted by $\mathcal{E}_{c}(x{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})$, satisfies the following condition:
@@
Proposition 3 shows that if there is an ellipsoid violation, then the point $m_{\lambda}$ belongs to the intersection Fig. 2(b):
@@
Although Algorithm 1 ensures safety according to Proposition 2, it does not guarantee target achievement. Agents within one-hop communication range may become immobilized at a certain position because any attempt to reach the target points could potentially violate the safety condition (7). We intentionally refrain from employing any enforcer to resolve deadlock scenarios. Instead, we allow the agents to learn how to avoid or recover from them through the training process.